OpenStudy (anonymous):

1. Factor. x2 − 6x + 9 (Points : 1) (x − 9)2 (x + 3)2 (x − 3)2 (x + 9)2 Question 2. 2. Which of the following is not a perfect square trinomial? (Points : 1) 81 + 18y + y2 144 + 24y + y2 121 + 11y + y2 100 − 20y + y2 Question 3. 3. Factor. m2 + 16m + 64 (Points : 1) Cannot be factored (m − 8)2 (m + 64)2 (m + 8)2

OpenStudy (chaise):

(a+b)^2=a^2+2ab+b^2 <--- this is a perfect square if (ab/2)^2=b (a-b)^2=a^2-2ab+b^2 (a-b)(a+b)=A^2+B^2 Does this help at all?

iYuko (iyuko):

I have a tutorial on factoring. Its a more advanced but if you would like..

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

2- 121 +11y+ Y^2 is not

OpenStudy (anonymous):

3 (m+8)^2

OpenStudy (anonymous):

1 (x-3)^2

OpenStudy (triciaal):

x2 − 6x + 9 what are the factors of +9 that add up to-6 -3, -3 (x-3)(x-3) which is (x-3)^2

OpenStudy (triciaal):

for 2) refer to the response by @chaise

OpenStudy (triciaal):

m2 + 16m + 64 what are the 2 factors of 64 that add up to 16 the signs are important 8 (m +8)(m + 8) = (m +8)^2